# Jiakang Yuan1,2, Xiangchao Yan2, Botian Shi2, Tao Chen1,🖂, Wanli Ouyang2 Bo Zhang2,‡,🖂, Lei Bai2,🖂, Yu Qiao2, Bowen Zhou2

### v2

@@ 6.1 Ideas Generation Process
             Compare $f_{s}$ with summary embeddings in $\mathbf{B}$ by $sim=f_{s}\cdot\mathbf{B}^{T}\in\mathbb{R}^{1\times len(\mathbf{B})}$.
@@
             end if
       end if
end for
Return: Independence list $\mathbf{R}$, len($\mathbf{R}$) == len($\mathbf{I}$)
@@
6.2 Experimental Verification Process
@@
7 More Details and Results of Experiments
@@
7.1 Details of Selected Tasks
7.1.1 Image Classification Task
Dataset: CIFAR-100. The CIFAR-100 dataset [19] includes colored natural images with a resolution of 32×32 pixels, categorized into 100 distinct classes. The training and test sets contain 50,000 and 10,000 images, respectively. We adopt a standard data augmentation scheme (i.e., RandomCrop, RandomHorizontalFlip, RandomRotation).
Implementation Details. We use WRN-28-10 [49] as our baseline. We partially refer to the codebase [^2] https://github.com/bmsookim/wide-resnet.pytorch. Our training process employs the SGD optimizer with the CosineAnuealing scheduler. The initial learning rate is set to 0.1, and we train the model for 200 epochs with a batch size of 128. We apply a weight decay of 5e-4 and a Nesterov momentum with a coefficient of 0.9.
7.1.2 3D Classification Task
Dataset: ModelNet40. ModelNet40 [45] is a synthetic object dataset that contains 12,311 3D CAD models covering 40 categories. The standard training/validation set of ModelNet40 carries 9843/2468 point clouds.
Implementation Details. We use PointNet [28] as our baseline. Following PointNet [28], we uniformly sample 1024 points on each object. We use the random scale, random dropout, and point shift during training and train the model for 200 epochs. The initial learning rate is set to 1e-3. We use Adam optimizer (weight decay=1e-4) and step learning rate decay (step size=20, gamma=0.7). Our implementation partially refers to codebase [^3] https://github.com/yanx27/Pointnet_Pointnet2_pytorch/tree/master.
7.1.3 Sentiment Classification Task
Dataset: SST-2. Stanford Sentiment Treebank (SST) [35] contains 11,855 one-sentence movie reviews extracted from Rotten Tomatoes. SST contains 215,154 unique manually labeled texts of varying lengths.
Implementation Details. Our code of sentiment classification tasks refers to the codebase [^4] https://github.com/YJiangcm/SST-2-sentiment-analysis. We fully fine-tune the BERT-base model for the classification task for 5 epochs with the learning rate 2e-5. The batch size is set to 32. We use the early stop mechanism during training (i.e., stop training if the accuracy of current epoch is lower than that of the previous epoch).
7.2 More Qualitative Results
@@
8 More analysis and Further Works
8.1 The Extensibility for the Research Task
Recently, lots of works have explored automatic idea generation [20, 34]. One limitation is that the novelty of an idea can only be judged through human scoring or large model scoring. However, in real scientific research, we need ideas that can lead to performance breakthroughs that can not be judged without experiments. Dolphin which includes ideas generation, experimental verification, and results feedback process can serve as an evaluation protocol. In the future, it can be combined with auto-idea generation works to assess the effectiveness of the idea generation.
8.2 Analysis on Future Works

### v3

@@ A.1 Ideas Generation Process
             Compare $f_{s}$ with summary embeddings in $\mathbf{B}$ by $sim=f_{s}\cdot\mathbf{B}^{T}\in\mathbf{R}^{1\times len(\mathbf{B})}$.
@@
return Independence list $\mathbf{R}$, len($\mathbf{R}$) == len($\mathbf{I}$)
@@
A.2 Experimental Verification Process
@@
Appendix B Additional Details and Results of Experiments
@@
B.1 Details of Selected Tasks
B.1.1 Image Classification Task
Dataset: CIFAR-100. The CIFAR-100 dataset Krizhevsky et al. (2009) includes colored natural images with a resolution of 32×32 pixels, categorized into 100 distinct classes. The training and test sets contain 50,000 and 10,000 images, respectively. We adopt a standard data augmentation scheme (i.e., RandomCrop, RandomHorizontalFlip, RandomRotation).
Implementation Details. We use WRN-28-10 Zagoruyko (2016) as our baseline. We partially refer to the codebase [^2] https://github.com/bmsookim/wide-resnet.pytorch. Our training process employs the SGD optimizer with the CosineAnuealing scheduler. The initial learning rate is set to 0.1, and we train the model for 200 epochs with a batch size of 128. We apply a weight decay of 5e-4 and a Nesterov momentum with a coefficient of 0.9.
B.1.2 3D Classification Task
Dataset: ModelNet40. ModelNet40 Wu et al. (2015) is a synthetic object dataset that contains 12,311 3D CAD models covering 40 categories. The standard training/validation set of ModelNet40 carries 9843/2468 point clouds.
Implementation Details. We use PointNet Qi et al. (2017a) as our baseline. Following PointNet Qi et al. (2017a), we uniformly sample 1024 points on each object. We use the random scale, random dropout, and point shift during training and train the model for 200 epochs. The initial learning rate is set to 1e-3. We use Adam optimizer (weight decay=1e-4) and step learning rate decay (step size=20, gamma=0.7). Our implementation partially refers to codebase [^3] https://github.com/yanx27/Pointnet_Pointnet2_pytorch/tree/master.
B.1.3 Sentiment Classification Task
Dataset: SST-2. Stanford Sentiment Treebank (SST) Socher et al. (2013) contains 11,855 one-sentence movie reviews extracted from Rotten Tomatoes. SST contains 215,154 unique manually labeled texts of varying lengths.
Implementation Details. Our code of sentiment classification tasks refers to the codebase [^4] https://github.com/YJiangcm/SST-2-sentiment-analysis. We fully fine-tune the BERT-base model for the classification task for 5 epochs with the learning rate 2e-5. The batch size is set to 32. We use the early stop mechanism during training (i.e.., stop training if the accuracy of current epoch is lower than that of the previous epoch).
B.1.4 MLE-bench
MLE-bench Chan et al. (2024) is designed to evaluate the capability of AI agents’ ML engineering. It is composed of 75 ML competitions from Kaggle. In our manuscript, we conduct experiment on 3 tasks in MLE-bench including detecting insults in social commentary challenge, tabular playground series dec 2021 challenge, and jigsaw toxic comment classification challenge.
B.2 Case Studies
@@
Appendix C Further Analysis and Future Work
C.1 The Extensibility for the Research Task
Recently, lots of works have explored automatic idea generation Li et al. (2024); Si et al. (2024). One limitation is that the novelty of an idea can only be judged through human scoring or large model scoring. However, in real scientific research, we need ideas that can lead to performance breakthroughs that can not be judged without experiments. Dolphin which includes ideas generation, experimental verification, and results feedback process can serve as an evaluation protocol. In the future, it can be combined with auto-idea generation works to assess the effectiveness of the idea generation.
C.2 Analysis on Future Works